OpenStudy (anonymous):

Solve x2 + 9x + 20 = 0. x = 2, x = 10 x = −2, x = −10 x = −5, x = −4 x = 5, x = 4

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (skullpatrol):

$$\huge x^2 + 9x + 20$$

OpenStudy (anonymous):

x=-5,-4

OpenStudy (anonymous):

You take get \(a\), \(b\), and \(c\). Then, to find \(x\): \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

So \[x = \frac {-9 \pm \sqrt {9^2 - 4 * 1 * 20}} {2 * 1}\]

OpenStudy (anonymous):

Is it A

OpenStudy (anonymous):

I gave you a formula with no variables. Just take a calculator and calculate it.

OpenStudy (anonymous):

Is it B

OpenStudy (anonymous):

Calculate it.

OpenStudy (anonymous):

I can tell you that it's C(x = -5, x = -4)

OpenStudy (anonymous):

@harliedanielfrye ?

OpenStudy (anonymous):

Yeah that i was gonna say

OpenStudy (anonymous):

Write the formula down please. That way you can just look at it when you need it.

OpenStudy (anonymous):

well my sister just toke so no need

OpenStudy (anonymous):

What does that even mean?